# Beyond RMSE and MAE: Introducing EAUC to unmask hidden bias and unfairness in dyadic regression models

### v1

@@ Abstract
Dyadic regression models, which predict real-valued outcomes for pairs of entities, are fundamental in many domains (e.g. predicting the rating of a user to a product in Recommender Systems) and promising and under exploration in many others (e.g. approximating the adequate dosage of a drug for a patient in personalized pharmacology). In this work, we demonstrate that non-uniformity in the observed value distributions of individual entities leads to severely biased predictions in state-of-the-art models, skewing predictions towards the average of observed past values for the entity and providing worse-than-random predictive power in eccentric yet equally important cases. We show that the usage of global error metrics like Root Mean Squared Error (RMSE) and Mean Absolute Error (MAE) is insufficient to capture this phenomenon, which we name eccentricity bias, and we introduce Eccentricity-Area Under the Curve (EAUC) as a new complementary metric that can quantify it in all studied models and datasets. We also prove the adequateness of EAUC by using naive de-biasing corrections to demonstrate that a lower model bias correlates with a lower EAUC and vice-versa. This work contributes a bias-aware evaluation of dyadic regression models to avoid potential unfairness and risks in critical real-world applications of such systems.
keywords:
Regression on Dyadic Data, Ethics in Artificial Intelligence, Fairness
PACS:
07.05.Mh
MSC:
[2020] 68T01
[…]
[inst1]organization=Universidade da Coruña, CITIC,addressline=Campus de Elviña s/n, postcode=15008, state=A Coruña, country=Spain
1 Introduction
@@
However, within these critical tasks, the presence of biases related to unfairness can have profound implications, such as disparate impacts on minority or vulnerable groups, unequal access to opportunities, and discriminatory decision-making processes 10.1145/3457607 10.1145/3494672 . From a legal perspective, regulations and guidelines are emerging globally to ensure fairness and ethics in AI systems. For instance, the European Union’s AI Act will regulate that AI systems must not be biased such they “perpetuate historical patterns of discrimination” euaiact , and the US government has emphasized that agencies and offices must act to “address algorithmic discrimination” whitehouse_2023 . Such regulations underscore the need for fairness in algorithmic decision-making to prevent discriminatory outcomes.
In the particular context of regression over dyadic data or dyadic regression, unfairness biases may manifest in real-world contexts as suboptimal recommendations that prioritize highly-rated items over more adequate ones for a user’s needs, or amplify harmful messages in social media, favouring the creation of toxic echo chambersnoordeh2020echo ge2020understanding .
@@
The rest of this work is structured as follows: Section 2 establishes the foundation by providing insights into dyadic regression and its history. Section 3 demonstrates the insufficiencies of the evaluation methodologies used in the State of the Art Following that, Section 4 presents our bias-aware metric EAUC, detailing how to calculate, use and interpret it to evaluate dyadic regression tasks.. Section 5 covers the datasets, models and implementation details of the experimental setup. Section 6 presents results and discussions, analyzing the prevalence of eccentricity bias and demonstrating the adequateness of EAUC to measure its severity in any model. Section 7 draws the conclusions of our study and outlines promising avenues for future research.
2 Background and Motivation
@@
2.1 Defining Dyadic Data and Dyadic Regression
@@
Beyond RS, regression over dyadic data can find applications in various critical domains like finance, to know the most profitable market price of a commodity for different buyers or countries vidmer2015prediction , or biomedicine, to accurately predict the adequate dosage of a drug required for an individual’s treatment ota2022application hu2018improvement . In these, unlike in L2R tasks, the precise magnitude of $\hat{r}_{ui}$ is crucial. As such, dyadic regression can affect people or their decisions, and predictions should be accurate, unbiased and fair due to their potential real-world impact.
2.2 State-of-the-Art Evaluation: Netflix Prize Competition and RMSE
For more than one decade, Root Mean Squared Error (RMSE) has stood as the prevailing evaluation metric for dyadic regression tasks, especially in the field of Recommender Systems. Its adoption can be attributed to seminal events, notably the Netflix Prize bennett2007netflix , a competition initiated by Netflix in 2006 that challenged researchers outperform the company’s existing movie recommendation system, Cinematch. This competition provided a fertile ground for innovation bell2007bellkor and brought significant attention to the field bell2007lessons hallinan2016recommended .
@@ 2.2 State-of-the-Art Evaluation: Netflix Prize Competition and RMSE
where $N$ is the total number of predictions, $r_{i}$ is the observed value, and $\hat{r_{i}}$ is the model’s prediction. In the Netflix Prize, observed values were discrete integer ratings $r_{ui}\in{1,2,3,4,5}$, and models predicted real-valued ratings $\hat{r}_{ui}\in[1,5]$.
RMSE gained prominence due to its intuitive interpretation (lower RMSE values indicating higher prediction accuracy), and the reliance on RMSE and other global error metrics like Mean Absolute Error (MAE) has continued to influence the field way after the Netflix Prize. Similar popular datasets like MovieLens harper2015movielens further cemented global error metrics as RMSE or MAE as the sole standard for evaluating regression tasks over dyadic data bobadilla2013recommender chen2017performance .
@@
3 Motivation: Limitations of RMSE in Dyadic regression tasks
@@
To exemplify the limitations of RMSE we turn to a practical example using the Netflix Prize data bennett2007netflix and a Matrix Factorization (MF) model rendle2020neural , a benchmark dataset and model from the State of the Art; both are further described in Sections 5.1 and 5.2, respectively.
Table 1 shows the Test RMSE for the samples of each unique rating (1 to 5) and the entire Test set. Although the system performs better for reviews with ratings 3 or 4, the most common in the data ($\sim 28\%$ and $\sim 33\%$ of ratings, respectively), predictions are always notably better than random, suggesting the absence of any important biases.
Table 1: RMSE values of an MF model in the Netflix Prize dataset, segregated by the rating (user’s score to a given movie) to be predicted.
@@
Figure 2: Eccentricity $Ecc_{ui}$ vs. prediction error $|\hat{r}_{ui}-r_{ui}|$ of an MF model and a uniform random predictor in the Netflix Prize dataset; mean and standard deviation of error over 5 runs is provided. The $y=x$ line corresponds to predicting always the $DMV_{ui}$ of each sample.
@@
4 Bias-aware Evaluation of Dyadic Regression Models
The preceding section unveiled the limitations of traditional metrics, like RMSE and MAE, in evaluating dyadic regression models. This section introduces EAUC, an alternative metric that can comprehensively detect and quantify eccentricity bias and its subsequent unfairness in model predictions.
Our objective is to propose an evaluation metric that provides a holistic view of dyadic regression models’ performance: a good model should excel not only in predicting near the $DMV_{ui}$ of the dyad, i.e. the expected value of the (u,i) interaction, but also in cases with higher eccentricity, where the model must prove it has understood $u$’s and $i$’s latent characteristics rather than memorizing $\bar{r}_{u}$ and $\bar{r}_{i}$. Such eccentric cases are equally essential, often representing critical edge cases of tasks where an AI system would be most useful.
In this work, we propose EAUC (Eccentricity-Area Under the Curve), a metric that provides a unified measure of predictive accuracy in both common and eccentric use cases. Our approach draws inspiration from area-based metrics like AUC-ROC, which assesses the overall discriminative power of classifiers.
4.1 Eccentricity-Area Under the Curve (EAUC)
@@
    For each test sample $j:(u,i,r_{ui},\hat{r}_{ui})$, calculate $(Ecc_{j},\epsilon_{j})$:
    • (a)
      $Ecc_{j}$ is the eccentricity of the observed value $r_{ui}$.
    • (b)
      $\epsilon_{j}$ is the absolute prediction error between $\hat{r}_{ui}$ and $r_{ui}$.
@@
    Integrate numerically to determine the area $A$ under the curve defined by eccentricities $Ecc$ and absolute errors $\epsilon$.
@@
1:Inputs: Train samples $D_{\text{train}}=\{(u,i,r_{ui})\}$ and test samples $D_{\text{test}}=\{(u,i,r_{ui},\hat{r}_{ui})\};u\in\mathcal{U},i\in\mathcal{I},r_{ui}\in\mathbb{R},\hat{r}_{ui}\in\mathbb{R}$
2:Outputs: Calculated score EAUC
3:Initialize $\bar{r}_{users}\leftarrow\mathbb{R}^{|\mathcal{U}|}$, $\bar{r}_{items}\leftarrow\mathbb{R}^{|\mathcal{I}|}$
4:for $j=1$ to $|\mathcal{U}|$ do
@@
6:end for
7:for $j=1$ to $|\mathcal{I}|$ do
8:     $\bar{r}_{items_{j}}\leftarrow\frac{1}{|D_{\text{train}_{j}}|}\sum_{(u,i,r_{ui})\in D_{\text{train}_{j}}}r_{ui}$
9:end for
@@
11:for $j=1$ to $|D_{\text{test}}|$ do
12:     $(u,i,r_{ui},\hat{r}_{ui})\leftarrow D_{\text{test}_{j}}$
13:     $Ecc_{j}\leftarrow|r_{ui}-\frac{1}{2}\cdot(\bar{r}_{users_{u}}+\bar{r}_{items_{i}})|$
14:     $\epsilon_{j}\leftarrow|\hat{r}_{ui}-r_{ui}|$
@@
16:Sort $(Ecc_{j},\epsilon_{j})$ pairs by ascending $Ecc_{j}$
17:Initialize $A\leftarrow 0$
18:for $j=2$ to $|D_{\text{test}}|$ do
19:     $\Delta A_{j}\leftarrow\frac{1}{2}\cdot(Ecc_{j}-Ecc_{j-1})\cdot(\epsilon_{j}+\epsilon_{j-1})$
@@
21:end for
22:EAUC $\leftarrow\frac{A}{(max(r_{ui})-min(r_{ui}))^{2}};(u,i,r_{ui},\hat{r}_{ui})\in D_{\text{test}}$
4.2 Interpretation and usage of EAUC
EAUC produces a value in the range [0, 1], where lower values indicate lower eccentricity bias and better predictive power. An ideal model with perfect predictions not biased towards $DMV_{ui}$ would achieve an EAUC of 0, and a naive predictor that always predicts the $DMV_{ui}$ of each test case would achieve an EAUC of 0.5, as long as it is tested in samples with eccentricities that span the range $[0,max(r_{ui})]$.
@@ 4.2 Interpretation and usage of EAUC
4.3 Measuring Dataset Difficulty for Eccentricity Bias
Using the same dyadic regression dataset, EAUC can be used to measure and compare the eccentricity bias of different models. Between datasets, however, a difference in EAUC could be caused not only by model adequateness, but also due to the characteristics or difficulty of the dataset. Dataset difficulty metrics are used in Machine Learning to assess the upfront difficulty of tasks or datasets and contextualize model performance li2014assessing ; collins2018evolutionary ; ho2002complexity . In dyadic data, dataset sparsity and scarcity are used to measure local and global quantities of information, but do not characterize the local distributions of observed values of individual entities, which may also influence the learning of models.
@@ 5.1 Datasets
In this work, we focus on dyadic regression datasets where each observation involves entities of two different types and an associated numerical value to be predicted. Table 2 shows basic statistics of the selected datasets, including the $D_{KS_{\mathcal{D}}}$ measurement of dataset difficulty. This selection provides a variety of data sources, context tasks, sample quantity and sparsity, and uniformity of local value distributions.
@@
Table 2: Basic statistics of the used datasets. Density ($|\mathcal{D}_{\text{train}}|/(|\mathcal{U}|\times|\mathcal{I}|)$) is indicative of the quantity of information per entity.
Dataset # Observations Density # Users # Items $\mathbf{D_{KS_{\mathcal{D}}}}$ Values
Movielens 100K [^1] https://grouplens.org/datasets/movielens/ harper2015movielens 100,000 6.30% 944 1,683 0.415 [1, 2, …, 5]
Movielens 1Mfootnotemark: harper2015movielens 1,000,209 4.50% 6,040 3,706 0.412 [1, 2, …, 5]
Movielens 10Mfootnotemark: harper2015movielens 10,000,054 1.30% 69,878 10,677 0.437 [0.5, 1, …, 4.5, 5]
Netflix Prize [^2] https://archive.org/download/nf_prize_dataset.tar bennett2007netflix 100,480,507 1.20% 480,189 17,770 0.429 [1, 2, …, 5]
Douban Monti [^3] https://github.com/fmonti/mgcnn monti2017geometric 136,066 1.50% 3,000 3,000 0.466 [1, 2, …, 5]
Tripadvisor pereznunez2021 491,170 0.20% 25,388 12,384 0.539 [1, 2, …, 5]
@@
Each of these datasets offers a scenario where the goal is to predict the numerical value a user would assign to a particular item (e.g., movies, TV series, restaurants). We apply a 90/10 random train/test data split, a common partitioning strategy that exposes models to the cold start problem meng2020exploring . As most users in the original Tripadvisor data only have 1 sample, we only keep those users and items with $\geq 10$ observed interactions. To maintain a strict dyadic regression protocol, no extra information per sample is available other than user and item identifiers.
5.2 Models

### v2

@@ Abstract
Dyadic regression models, which predict real-valued outcomes for pairs of entities, are fundamental in many domains (e.g. predicting the rating of a user to a product in Recommender Systems) and promising and under exploration in others (e.g. approximating the adequate dosage of a drug for a patient in personalized pharmacology). In this work, we prove that non-uniformity in the observed value distributions of individual entities leads to severely biased predictions in state-of-the-art models, skewing predictions towards the average of observed past values for the entity and providing worse-than-random predictive power in eccentric yet equally important cases. We show that global error metrics like Root Mean Squared Error (RMSE) and Mean Absolute Error (MAE) are insufficient to capture this phenomenon, which we name eccentricity bias, and we introduce Eccentricity-Area Under the Curve (EAUC) as a novel complementary metric that can quantify it in all studied models and datasets. We also show the adequateness of EAUC, demonstrating with naive de-biasing corrections that a lower model bias correlates with a lower EAUC and vice-versa. This work contributes a bias-aware evaluation of dyadic regression models to prevent unfairness and risks in critical real-world applications of such systems.
Index Terms:
Regression on Dyadic Data, Fairness, Ethics in Artificial Intelligence, Machine Learning
I Introduction
@@
However, within these critical tasks, the presence of biases related to unfairness can have profound implications, such as disparate impacts on minority or vulnerable groups, unequal access to opportunities, and discriminatory decision-making processes [1] [2]. From a legal perspective, regulations and guidelines are emerging globally to ensure fairness and ethics in AI systems. For instance, the European Union’s AI Act will regulate that AI systems must not be biased such they “perpetuate historical patterns of discrimination” [3], and the US government has emphasized that agencies and offices must act to “address algorithmic discrimination” [4]. Such regulations underscore the need for fairness in algorithmic decision-making to prevent discriminatory outcomes.
In the particular context of regression over dyadic data or dyadic regression, unfairness biases may manifest in real-world contexts as suboptimal recommendations that prioritize highly-rated items over more adequate ones for a user’s needs, or amplify harmful messages in social media, favouring the creation of toxic echo chambers[5][6].
@@
II Background and Motivation
@@
II-A Defining Dyadic Data and Dyadic Regression
@@
Beyond RS, regression on dyadic data finds applications in various critical domains like finance, to know the most profitable market price of a commodity for different buyers or countries [7], or biomedicine, to predict the adequate dosage of a drug required for an individual’s treatment [8][9]. In these, unlike in L2R tasks, the precise magnitude of $\hat{r}_{ui}$ is crucial. As such, dyadic regression can affect people or their decisions, and predictions should be accurate, unbiased and fair due to their potential real-world impact.
II-B State-of-the-Art Evaluation: Netflix Prize Competition and RMSE
For more than one decade, Root Mean Squared Error (RMSE) has stood as the prevailing evaluation metric for dyadic regression tasks, especially in the field of Recommender Systems. Its adoption can be attributed to seminal events, notably the Netflix Prize [10], a competition initiated by Netflix in 2006 that challenged researchers outperform the company’s existing movie recommendation system, Cinematch. This competition provided a fertile ground for innovation [11] and brought significant attention to the field [12][13].
@@ II-B State-of-the-Art Evaluation: Netflix Prize Competition and RMSE
where $N$ is the total number of predictions, $r_{i}$ is the observed value, and $\hat{r_{i}}$ is the model’s prediction. In the Netflix Prize, observed values were discrete ratings $r_{ui}\in{1,2,3,4,5}$, and models predicted real-valued ratings $\hat{r}_{ui}\in[1,5]$.
RMSE gained prominence due to its intuitive interpretation (lower RMSE values indicating higher prediction accuracy), and the reliance on RMSE and other global error metrics like Mean Absolute Error (MAE) has continued to influence the field way after the Netflix Prize. Similar popular datasets like MovieLens [14] further cemented global error metrics as RMSE or MAE as the sole standard for evaluating regression tasks over dyadic data [15][16].
@@
III Motivation: Limitations of RMSE in Dyadic regression tasks
@@
To exemplify the limitations of RMSE we turn to a practical example using the Netflix Prize data [10] and a Matrix Factorization (MF) model [17], a benchmark dataset and model from the State of the Art; both are further described in Sections V-A and V-B, respectively.
Table I shows the Test RMSE for the samples of each unique rating (1 to 5) and the entire Test set. Although the system performs better for reviews with ratings 3 or 4, the most common in the data ($\sim 28\%$ and $\sim 33\%$ of ratings, respectively), predictions are always notably better than random, suggesting the absence of any important biases.
TABLE I: RMSE of an MF model, seggregated by rating
@@
  • Values obtained by an MF model in the Netflix Prize dataset. The rating is the value $r_{i}$ to be predicted, i.e. a user’s score to a given movie.
@@
Figure 2: Eccentricity $Ecc_{ui}$ vs. prediction error $|\hat{r}_{ui}-r_{ui}|$ of an MF model and a uniform random predictor in the Netflix dataset; mean and standard deviation of error over 5 runs is provided. The $y=x$ line corresponds to predicting always the $DMV_{ui}$ of each sample.
@@
IV Bias-aware Evaluation of Dyadic Regression Models
This section introduces EAUC, an alternative metric that can detect and quantify eccentricity bias and its subsequent unfairness in dyadic regression models’ predictions by providing a holistic view of their performance: a good model should excel not only in predicting near the $DMV_{ui}$ of the dyad, i.e. the expected value of the (u,i) interaction, but also in cases with higher eccentricity, where the model must prove it has understood $u$’s and $i$’s latent characteristics rather than memorizing $\bar{r}_{u}$ and $\bar{r}_{i}$. Such eccentric cases are equally essential, often representing critical edge cases of tasks where an AI system would be most useful.
IV-A Eccentricity-Area Under the Curve (EAUC)
@@
    For each test sample $j:(u,i,r_{ui},\hat{r}_{ui})$, calculate:
    • $Ecc_{j}$, the eccentricity of the observed value $r_{ui}$.
    • $\epsilon_{j}$, the absolute prediction error $|\hat{r}_{ui}-r_{ui}|$.
@@
    Integrate to determine the area $A$ under the curve defined by eccentricities $Ecc$ and absolute errors $\epsilon$.
@@
Inputs:
  • Train samples $D_{\text{train}}=\{(u,i,r_{ui})\}$
  • Test samples $D_{\text{test}}=\{(u,i,r_{ui}$
  • ($u\in\mathcal{U},i\in\mathcal{I},r_{ui}\in\mathbb{R},\hat{r}_{ui}\in\mathbb{R}$)
Outputs: Calculated score EAUC
Initialize $\bar{r}_{users}\leftarrow\mathbb{R}^{|\mathcal{U}|}$, $\bar{r}_{items}\leftarrow\mathbb{R}^{|\mathcal{I}|}$
for $j=1$ to $|\mathcal{U}|$ do
@@
end for
for $j=1$ to $|\mathcal{I}|$ do
     $\bar{r}_{items_{j}}\leftarrow\frac{1}{|D_{\text{train}_{j}}|}\sum_{(u,i,r_{ui})\in D_{\text{train}_{j}}}r_{ui}$
end for
@@
for $j=1$ to $|D_{\text{test}}|$ do
     $(u,i,r_{ui},\hat{r}_{ui})\leftarrow D_{\text{test}_{j}}$
     $Ecc_{j}\leftarrow|r_{ui}-\frac{1}{2}\cdot(\bar{r}_{users_{u}}+\bar{r}_{items_{i}})|$
     $\epsilon_{j}\leftarrow|\hat{r}_{ui}-r_{ui}|$
@@
Sort $(Ecc_{j},\epsilon_{j})$ pairs by ascending $Ecc_{j}$
Initialize $A\leftarrow 0$
for $j=2$ to $|D_{\text{test}}|$ do
     $\Delta A_{j}\leftarrow\frac{1}{2}\cdot(Ecc_{j}-Ecc_{j-1})\cdot(\epsilon_{j}+\epsilon_{j-1})$
@@
end for
EAUC $\leftarrow\frac{A}{(max(r_{ui})-min(r_{ui}))^{2}};(u,i,r_{ui},\hat{r}_{ui})\in D_{\text{test}}$
IV-B Interpretation and usage of EAUC
EAUC produces a value in the range [0, 1], where lower values indicate lower eccentricity bias and better predictive power. An ideal model with perfect predictions not biased towards $DMV_{ui}$ would achieve an EAUC of 0, and a naive predictor that always predicts the $DMV_{ui}$ would achieve an EAUC of 0.5, as long as it is tested in samples with eccentricities that span the range $[0,max(r_{ui})]$.
@@ IV-B Interpretation and usage of EAUC
IV-C Measuring Dataset Difficulty for Eccentricity Bias
Using the same dyadic regression dataset, EAUC can be used to measure and compare the eccentricity bias of different models. Between datasets, however, a difference in EAUC could be caused not only by model adequateness, but also due to the characteristics or difficulty of the dataset. Dataset difficulty metrics are used in Machine Learning to assess the upfront difficulty of tasks or datasets and contextualize model performance [18, 19, 20]. In dyadic data, dataset sparsity and scarcity are used to measure local and global quantities of information, but do not characterize the local distributions of observed values of individual entities, which may also influence the learning of models.
@@ V-A Datasets
We focus on dyadic regression datasets where each observation involves entities of two different types and an associated numerical value to be predicted. Table II shows basic statistics of the selected datasets, including the $D_{KS_{\mathcal{D}}}$ measurement of dataset difficulty. This selection provides a variety of data sources, context tasks, sample quantity and sparsity, and uniformity of local value distributions.
TABLE II: Basic statistics of the used datasets.
Dataset Observations Density Users Items $\mathbf{D_{KS_{\mathcal{D}}}}$
Movielens 100K [^1] https://grouplens.org/datasets/movielens/ [14] 100,000 6.30% 944 1,683 0.415
Movielens 1Mfootnotemark: [14] 1,000,209 4.50% 6,040 3,706 0.412
Movielens 10Mfootnotemark: [14] 10,000,054 1.30% 69,878 10,677 0.437
Netflix Prize [^2] https://archive.org/download/nf_prize_dataset.tar [10] 100,480,507 1.20% 480,189 17,770 0.429
Douban Monti [^3] https://github.com/fmonti/mgcnn [21] 136,066 1.50% 3,000 3,000 0.466
Tripadvisor [22] 491,170 0.20% 25,388 12,384 0.539
@@
  • The density ($|\mathcal{D}_{\text{train}}|/(|\mathcal{U}|\times|\mathcal{I}|)$) indicates the quantity of implicit feedback information available per entity (user or item).
@@
Each of these datasets offers a scenario where the goal is to predict the numerical value a user would assign to a particular item (e.g., movies, TV series, restaurants). We apply a 90/10 random train/test data split, a common partitioning strategy that exposes models to the cold start problem [23]. As most users in the original Tripadvisor data only have 1 sample, we only keep those users and items with $\geq 10$ observed interactions. To maintain a strict dyadic regression protocol, no extra information per sample is available other than user and item identifiers.
V-B Models